the difference of two numbers is 15. five times the smaller is the same as 9 less than twice the larger. find the numbers

Answer by mananth(16949)

one number is x
other number is x+15
..
5x=(2(x+15)-9)
5x=2x+30-9
5x-2x=21
3x=21
/3
x= 7
7+15 = 22 is the second number
..
7 & 22 are the numbers
..
m.ananth@hotmail.ca